• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name jpchen

    Log entry time 00:05:20 on August 9,2001

    Entry number 70025

    keyword=swing shift summary

    8/8 swing shift summary
    Shift workers: J. P. Chen, K. McCormick, O. Gayou.

    16:00
    Shift starts with smooth data taking at E=3.465 GeV, theta=18.6 degree, P=1.6 GeV.
    Beam half wave plate in, target at 90 degree, with target half wave plate in.

    20:55-21:25 NMR/EPR measurements after run 2705/21705
    NMR=35.2%, EPR=37.8%

    22:50 MCC called that they have communication problems, beam is
    off and no estimation until AES look into the problem.
    23:15 Beam back.
    23:20 Run 2708/21708 were short runs due to problem with Parity
    DAQ run 1839 could not start taking events.
    23:25 After kcoda for the parity DAQ, restarts fine.

    Left run# Good e right good e Total (previous=4.146M)
    2699 407K 21699 394K 4.947M
    2700 388K 21700 376K 5.711M
    2701 374K 21701 362K 6.447M
    2702 345K 21702 388K 7.230M
    2703 297K 21703 286K 7.813M
    2704 343K 21704 332K 8.488M
    2705 400K 21705 391K 9.279M
    2706 395K 21706 383K 10.057M
    2707 382K 21707 369K 10.808M
    2708/21708 are short due to parity daq crash
    2709/21709 still are being analyzed at the end of this shift